Let E and F be two events of an experiment with sample space S. Suppose P(E) = 0.2, P(F) = 0.1 and P(E n F) = 0.05. Compute.

a.      P(E u F)

b.     P(Ec n Fc)

c.      P(E n F)



1
Expert's answer
2021-12-02T18:54:17-0500

a.


"P(E\cup F)=P(E)+P(F)-P(E\cap F)"

"P(E\cup F)=0.2+0.1-0.05=0.25"

b.


"P(E^C\cap F^C)=1-P(E\cup F)"

"P(E^C\cap F^C)=1-0.25=0.75"

c.


"P(E^C\cup F^C)=P(E^C)+P(F^C)-P(E^C\cap F^C)"

"=1-P(E)+1-P(F)-P(E^C\cap F^C)"

"P(E^C\cup F^C)=1-0.2+1-0.1-0.75=0.95"

"P(E^C\cup F^C)=1-P(E\cap F)"

"P(E^C\cup F^C)=1-0.05=0.95"